# Intermediate Relation Size Bounds for Select-Project-Join Query Plans: Asymptotically Tight Characterizations

### v1

@@ 3. Intermediate Degree and Main Theorem Statement
Let $\Sigma$ be a set of keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$, and let $d\in\mathbb{Q}$. The plan $p$ has output degree $\leq d$ if there exists a function $h:\mathbb{N}\to\mathbb{N}$, with $h(M)\in O(M^{d})$, such that for every structure $\mathbf{D}$ satisfying $\Sigma$, it holds that $|\mathsf{out}(p,\mathbf{D})|\leq h(M_{\mathbf{D}})$. We say that $p$ has output degree $d$ if it has output degree $\leq d$, but for all $\epsilon>0$, it does not have output degree $\leq(d-\epsilon)$.
@@
Let $\Sigma$ be a set of keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$, and let $d\in\mathbb{Q}$. The plan $p$ has intermediate degree $\leq d$ if there exists a function $h:\mathbb{N}\to\mathbb{N}$, with $h(M)\in O(M^{d})$, such that for every structure $\mathbf{D}$ satisfying $\Sigma$, and for every subplan $q$ of $p$, it holds that $|\mathsf{out}(q,\mathbf{D})|\leq h(M_{\mathbf{D}})$. We say that $p$ has intermediate degree $d$ if it has intermediate degree $\leq d$, but for all $\epsilon>0$, it does not have intermediate degree $\leq(d-\epsilon)$.
Note that the above definitions of output degree and intermediate degree are with respect to the keys $\Sigma$, but we omit $\Sigma$ when it is clear from context for the sake of readability. Given a set of constraints $\Sigma$, we say that an SPJ plan $p$ has the best possible intermediate degree if there exists some $d\in\mathbb{Q}$ such that $p$ has intermediate degree $d$, and for every SPJ plan $p^{\prime}$ that is $\Sigma$-semantically equivalent to $p$, and there is no value $\epsilon>0$ such that the plan $p^{\prime}$ has intermediate degree $\leq(d-\epsilon)$. In other words, no $\Sigma$-semantically equivalent plan achieves an intermediate degree that is strictly lower than that of $p$.
We analyze a plan $p$ by means of its so called $p$-representation, an open structure $(\mathbf{A},\overline{a})$ for which it holds that, for any structure $\mathbf{D}$, we have that $\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$. A formal treatment of $p$-representation and related notions utilized to show our results will follow in Section 4. We are now ready to formulate our main theorem.
@@
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ be a $p$-representation. Let $(\mathbf{C},\overline{c})$ be a core of $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$; then there exists a well-behaved plan $p^{\prime}$ where
  • $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$,
  • $p^{\prime}$ has intermediate degree $d$, and
  • $p^{\prime}$ has the best possible intermediate degree.
There exists an exponential-time algorithm that, given an SPJ plan $p$, outputs such a plan $p^{\prime}$; the plan $p^{\prime}$ has size polynomial in the size of $p$.
We remark that given an SPJ plan $p$ the value $d$ is computable in exponential time; this follows quite directly from the polynomial time computability of the color number, which follows from results in Sections 3 and 4 from (Gottlob et al., 2012). The algorithm provided by Theorem 3.3 passes from a plan $p$ to a second plan $p^{\prime}$ with the best possible intermediate degree. A natural question that arises is how much time it takes to evaluate $p^{\prime}$ on a given structure $\mathbf{D}$. The following proposition, which can be argued elementarily, gives an answer to this question: given a well-behaved plan, apart from a multiplicative overhead depending on the plan, the time needed is at most $|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d}$, where $d$ is the intermediate degree of the plan—so in essence, the intermediate degree of the plan dictates the polynomial degree of evaluation.
@@
There exists an algorithm that, given a well-behaved SPJ plan $p$ of intermediate degree $\leq d$ and a structure $\mathbf{D}$, evaluates $p$ on $\mathbf{D}$ within time $O(|p|^{2}*|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d})$.
@@
To argue this proposition, we consider an algorithm that evaluates the plan $p$ inductively. It suffices to show that each subplan $q\in\mathsf{subplans}(p)$ can be computed in time $O(|p|*|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d})$. By the definition of intermediate degree, each subplan $q$ has at most $O(|M_{\mathbf{D}}|^{d})$ answers (tuples in $\mathsf{out}(p,\mathbf{D})$). When $q$ is a basic plan, the required bound holds immediately. When $q$ is a select or project subplan, i.e., of the form $\sigma_{\theta}(q^{\prime})$ or $\pi_{j_{1},\ldots,j_{n}}(q^{\prime})$, respectively, the answers of $q$ can be naturally derived from the answers of the immediate subplan $q^{\prime}$, within the given time bound. The key point is that, when $q$ is a join subplan, by well-behavedness, it is of the form $\Join_{\theta}(q_{1},\ldots,q_{\ell})$ and $\mathsf{out}(q,\mathbf{D})$ can be computed by maintaining a relation that is a subset of $\mathsf{out}(q_{1},\mathbf{D})\times\mathsf{dom}(\mathbf{D})$. When evaluating the multiway join, we can thus evaluate the sequence of binary joins $\Join_{\theta_{\ell-1}}(\ldots(\Join_{\theta_{2}}(\Join_{\theta_{1}}(q_{1},q_{2}),q_{3}),\ldots),q_{\ell})$, where $\theta_{1},\ldots,\theta_{\ell-1}$ are the corresponding equalities from $\theta$ for the join partners in question. Each binary join can be evaluated in time $O(|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d})$ by (Flum et al., 2002), and in total, there are at most $|p|$ many joins, giving us the desired runtime of $O(|p|*|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d})$ to evaluate the subplan $q$.
Note that our main theorem only considers constraints which are unary keys. This is a limitation that arises from the literature regarding the bounds on the size of the output of a plan and is challenging to overcome. To extend the bounds we inherit from the literature, one is likely to make significant progress towards a long-standing open problem in information theory (Gottlob et al., 2012).
Proving Theorem 3.3 essentially requires us to show two statements: (i) we can synthesize a plan, which is $\Sigma$-equivalent to the original one and has intermediate degree $\leq d$ and (ii) any plan which is $\Sigma$-equivalent to the original one has a subplan with output degree at least $d$. We will formalize and prove (i) and (ii) in Theorems 5.1 and 5.2, respectively, in Section 5.
@@
To show the theorem, we make two more basic observations on the introduced degree notions. We begin by presenting a definition that will aid us in arguing that a plan has a certain output degree. We say that $p$ has output degree $\geq d$ if there exists a function $g:\mathbb{N}\to\mathbb{N}$ with $g(M)\in\Omega(M^{d})$ and there exists an infinite sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures with $M_{\mathbf{D}_{n}}$ unbounded such that (for all $n\geq 1$) $|\mathsf{out}(p,\mathbf{D}_{n})|\geq g(M_{\mathbf{D}_{n}})$. We then have the following lemma.
@@
Theorem 3.3 can then be argued as follows: Given a plan $p$, we first compute the open structure $(\mathbf{A},\overline{a})$, which can be done in polynomial time by Theorem 4.4. Then, the algorithm of Theorem 3.3 computes $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, again in polynomial time. It then computes the core of this open structure; this can be performed in exponential time. Let $(\mathbf{C},\overline{c})$ be this core. Note that the size of $(\mathbf{C},\overline{c})$ is bounded above by the size of $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, since the core of a structure is a substructure of the structure. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. From $(\mathbf{C},\overline{c})$, we can then synthesize the desired plan $p^{\prime}$ of intermediate degree $\leq d$ by statement (i), formalized in Theorem 5.1, from above. Consider any plan $p^{\prime\prime}$ that is $\Sigma$-semantically equivalent to $p^{\prime}$. By statement (ii), formalized in Theorem 5.2, $p^{\prime\prime}$ has a subplan with output degree $\geq d$. For any $\epsilon>0$, by Lemma 3.5, this subplan does not have output degree $\leq(d-\epsilon)$, and so by Lemma 3.6, $p^{\prime\prime}$ does not have intermediate degree $\leq(d-\epsilon)$. Hence, we have computed a $\Sigma$-equivalent plan $p^{\prime}$ of polynomial size and of intermediate degree $d$, which is the best possible intermediate degree.
In the next section, we will introduce the connection from plans to open structures and auxiliary notions, which are needed to prove the theorems formalizing statements (i) and (ii) in Section 5.
4. Relating queries to structures
There is a known correspondence between Boolean SPJ plans and structures (Chandra and Merlin, 1977); by a Boolean SPJ plan, we mean one that evaluates to true or false (or alternatively, a relation of arity 00) on a given structure $\mathbf{D}$. This correspondence yields that each such plan can be translated to a structure $\mathbf{A}$ such that, for an arbitrary structure $\mathbf{D}$, the plan is true on $\mathbf{D}$ if and only if there is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$. We prove a version of this correspondence in Theorem 4.4 for SPJ plans that can output relations of any finite arity, and not just relations of arity 00. In order to formulate our version, we utilize the notion of open structure.
@@ 4. Relating queries to structures
Apart from showing that any SPJ plan $p$ can be translated to a $p$-representation, we also want to provide a tree decomposition of this open structure with certain additional properties, which we call $p$-decomposition. This allows us to relate plans to structural, decomposition-based measures. While the translation that we present is, on a high level, along the lines of known translations from SPJ plans to conjunctive queries, we emphasize that our translation also provides a tree decomposition that will allow us to analyze the intermediate degree. In particular, for a given plan, each subplan will correspond to a vertex of the tree decomposition whose associated set of elements corresponds to the free variables of the subplan.
We say a tree decomposition $(T,\chi)$ is rooted if there is a distinguished node $v_{0}\in V(T)$, called the root of $T$. The formal definition of $p$-decomposition follows.
@@
Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ a $p$-representation. A $p$-decomposition of $(\mathbf{A},\overline{a})$ is a rooted tree decomposition $E=(T,\chi)$ of $(\mathbf{A},\overline{a})$ such that there exists
  • a bijection $\alpha:\mathsf{subplans}(p)\to V(T)$, where $\alpha(p)$ is the root of $T$, and
  • a map $\beta:\mathsf{subplans}(p)\to\mathsf{dom}(\mathbf{A})^{*}$, where $\beta(p)$ is equal to $\overline{a}$,
where for each $q\in\mathsf{subplans}(p)$, we have that $\{\beta(q)\}=\chi(\alpha(q))$.
In the definition of $p$-decomposition, each subplan $q$ of $p$ corresponds to a vertex of the tree $T$, via $\alpha$, and for each such subplan $q$, it holds that $\beta(q)$ is a tuple that lists, in some order (and possibly with repetition), the elements of the set to which $\chi$ maps q’s tree vertex.
Let $p$ be an SPJ plan over signature $\sigma$, $(\mathbf{A},\overline{a})$ be a $p$-representation, and $E=(T,\chi)$ be a rooted $p$-decomposition $E=(T,\chi)$ of $(\mathbf{A},\overline{a})$ with associated mappings $\alpha$ and $\beta$. For a vertex $v\in T$, let $S_{v}$ be the union of $\chi(u)$ over all descendants $u$ of $v$ (including $v$ itself) and let $\mathbf{A}_{\leq v,E}$ denote the induced substructure $\mathbf{A}[S_{v}]$. We say a subplan $q\in\mathsf{subplans}(p)$ satisfies the containment property w.r.t. a $p$-decomposition if, for each structure $\mathbf{D}$ over $\sigma$, the relationship $\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}_{\leq\alpha(q),E},\beta(q),\mathbf{D})$ holds. Further, we say a $p$-decomposition satisfies the containment property, if every subplan $q\in\mathsf{subplans}(p)$ satisfies the containment property w.r.t. the $p$-decomposition. The containment property will allow us to lower bound the sizes of intermediate relations of $p$ using a $p$-representation along with a $p$-decomposition satisfying the containment property.
@@
Consider the signature $\sigma$ containing the binary relations $R/2$ and $S/2$ and the SPJ plan $p=\Join_{(1=3)}(R,\pi_{1}(S))$ over $\sigma$. Then it can be readily seen, with structure $\mathbf{A}=\{R(a,b),S(a,d)\}$ and tuple $\overline{a}=(a,b,a)$, that $(\mathbf{A},\overline{a})$ is a $p$-representation. Consider the following trees:
On the left, we have the plan $p$ represented as a tree (each node represents the subplan of its subtree), while on the right we have some tree $T$. We provide mappings $\alpha$, $\beta$, and $\chi$ such that $(T,\chi)$ is a $p$-decomposition via mappings $\alpha$ and $\beta$. First, $\alpha$ is the obvious bijection from the subplans of $p$ to $V(T)$ with $\alpha(p)=v_{0}$, $\alpha(R)=v_{1}$, $\alpha(\pi_{1}(S))=v_{2}$, and $\alpha(S)=v_{3}$. Further, we define $\beta$ as $\beta(p)=\overline{a}$, $\beta(R)=(a,b)$, $\beta(\pi_{1}(S))=(a)$, and $\beta(S)=(a,d)$. Lastly, define $\chi(\alpha(q))$ as $\{\beta(q)\}$ for all subplans $q$ of $p$. It can be readily verified that $(T,\chi)$ is indeed a $p$-decomposition, and that it satisfies the containment property.
We show with the following theorem, that a $p$-representation and a $p$-decomposition with the above properties always exist and are efficiently computable.

### v2

@@ 3.1 Degree notions
Let $\Sigma$ be a set of keys over signature $\sigma$. Let $p$ be a plan over $\sigma$, and let $d\in\mathbb{Q}$. The plan $p$ has output degree $\leq d$ if there exists a function $h:\mathbb{N}\to\mathbb{N}$, with $h(M)\in O(M^{d})$, such that for every structure $\mathbf{D}$ satisfying $\Sigma$, it holds that $|\mathsf{out}(p,\mathbf{D})|\leq h(M_{\mathbf{D}})$. We say that $p$ has output degree $d$ if it has output degree $\leq d$, but for all $\epsilon>0$, it does not have output degree $\leq(d-\epsilon)$.
@@
Let $\Sigma$ be a set of keys over signature $\sigma$. Let $p$ be a plan over $\sigma$, and let $d\in\mathbb{Q}$. The plan $p$ has intermediate degree $\leq d$ if there exists a function $h:\mathbb{N}\to\mathbb{N}$, with $h(M)\in O(M^{d})$, such that for every structure $\mathbf{D}$ satisfying $\Sigma$, and for every subplan $q$ of $p$, it holds that $|\mathsf{out}(q,\mathbf{D})|\leq h(M_{\mathbf{D}})$. We say that $p$ has intermediate degree $d$ if it has intermediate degree $\leq d$, but for all $\epsilon>0$, it does not have intermediate degree $\leq(d-\epsilon)$.
Note that the above definitions of output degree and intermediate degree are with respect to the keys $\Sigma$, but we omit $\Sigma$ when it is clear from context for the sake of readability. Let $\Sigma$ be a set of keys. We say that a plan $p$ has the SPJ-best possible intermediate degree if there exists $d\in\mathbb{Q}$ such that $p$ has intermediate degree $d$, and for every SPJ plan $p^{\prime}$ that is $\Sigma$-semantically equivalent to $p$, there is no value $\epsilon>0$ such that the plan $p^{\prime}$ has intermediate degree $\leq(d-\epsilon)$. In other words, no $\Sigma$-semantically equivalent SPJ plan achieves an intermediate degree that is strictly lower than that of $p$. Analogously, we say that a plan $p$ has the SPJU-best possible intermediate degree if there exists $d\in\mathbb{Q}$ such that $p$ has intermediate degree $d$, and for every SPJU plan $p^{\prime}$ that is $\Sigma$-semantically equivalent to $p$, there is no value $\epsilon>0$ such that the plan $p^{\prime}$ has intermediate degree $\leq(d-\epsilon)$. In the appendix, we give a number of examples of plans along with computations of the output degree and the best possible intermediate degree.
3.2 Main theorem statement
We are now ready to formulate our main theorem.
@@ 3.2 Main theorem statement
There exists an exponential-time algorithm that, given a signature $\sigma$, a set $\Sigma$ of unary keys over $\sigma$, and an SPJU plan $p$ over $\sigma$, outputs a SPJU plan $p^{\prime}$ satisfying the following properties: $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$, $p^{\prime}$ has the SPJU-best possible intermediate degree.
By an exponential-time algorithm, we mean an algorithm that runs within time $O(2^{Q(n)})$, where $n$ is the input size and $Q$ is a polynomial. The principal argumentation needed to establish our main theorem is located in the case where this theorem is specialized to SPJ queries. In this case, which we present next, we can describe the intermediate degree of the output plan quite directly, but require the following notion to do so. A $p$-representation of a plan $p$ is an open structure $(\mathbf{A},\overline{a})$ where, for any structure $\mathbf{D}$, we have $\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$. A formal treatment of $p$-representation and related notions utilized to show our results will follow in Subsection 4.1.
Theorem 3.4.
@@
There exists an exponential-time algorithm that, given a signature $\sigma$, a set $\Sigma$ of unary keys over $\sigma$, and an SPJ plan $p$ over $\sigma$, outputs a SPJ plan $p^{\prime}$ satisfying the following properties: $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$, $p^{\prime}$ has the SPJ-best possible intermediate degree. Moreover, suppose that $p$ is an SPJ plan over $\sigma$ and that $(\mathbf{A},\overline{a})$ is a $p$-representation; let $(\mathbf{C},\overline{c})$ be a core of $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, and let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. Then, when the algorithm is given $p$ as input, the output plan $p^{\prime}$ has intermediate degree $d$, and also has size polynomial in the size of $p$.
Some remarks are in order. First, note that in Theorem 3.4, the intermediate degree of each plan $p^{\prime}$ can be described quite explicitly. Note also that given an SPJ plan $p$, the intermediate degree $d$ of the plan $p^{\prime}$ given by the algorithm of Theorem 3.4 is computable in exponential time from $p$; this follows directly from the exponential-time computability of $(\mathbf{C},\overline{c})$ from $p$ and the polynomial time computability of the color number, which follows from results in Sections 3 and 4 from [10]. In turn, given an SPJU plan $p$, the intermediate degree of a plan $p^{\prime}$ output by Theorem 3.3’s algorithm is also computable in exponential time from $p$, which follows from the above fact and the construction of $p^{\prime}$ in the proof of Theorem 3.3. We also wish to remark that, while our main theorem only considers constraints which are unary keys, this limitation arises from the literature on output degree bounds; extending the literature’s bounds is likely to be challenging, for this would constitute significant progress towards a long-standing open problem in information theory [10, Section 1].
3.3 Algorithmic aspects
We here discuss algorithms for evaluating well-behaved plans in time that is close to a plan’s intermediate degree; this issue was motivated in the introduction. We define a plan $p$ to be well-behaved if for each subplan $q$ that is a join $\Join_{\theta}(p_{1},\ldots,p_{\ell})$ of plans $p_{1},\ldots,p_{\ell}$ having arities $m_{1},\ldots,m_{\ell}$, the following is satisfied, when we let $s$ denote the sum $m_{1}+\cdots+m_{\ell}$: there exists an index $k\in[s]$ such that for each index $i\in[s]\setminus([m_{1}]\cup\{k\})$, there exists an index $j\in[m_{1}]\cup\{k\}$ such that $(i=j)\in\theta$. In words, each join adds at most one column of new information to the plan $p_{1}$: letting $J$ denote the set containing (1) the indices in $m_{1}]$, which correspond to $p_{1}$, and (2) another chosen index $k$; it must be that every index $i\notin J$ is equated with an index in $J$. This property allows for the efficient computation of such joins and, thus, well-behaved plans; we will make this formal in Proposition 3.6. First, we point out that our minimization algorithms can be implemented in such a way that they always output well-behaved plans.
Theorem 3.5.
The algorithms of Theorems 3.3 and 3.4 can be implemented in such a way so that each plan $p^{\prime}$ output by them is well-behaved.
This easily follows from the fact that the algorithm of Theorem 4.7 always outputs well-behaved SPJ plans and that the SPJU plan output by the algorithm of Theorem 3.3 is simply the union of well-behaved SPJ plans, which is also a well-behaved plan. The following proposition shows that, given a well-behaved plan, apart from a multiplicative overhead depending on the plan, the time needed to algorithmically evaluate the plan on a database $\mathbf{D}$ is at most $|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d}$, where $d$ is the intermediate degree of the plan—so in essence, the intermediate degree of the plan dictates the polynomial degree of evaluation (as a function of $M_{\mathbf{D}}$). To argue this, we in essence use the argumentation for [2, Theorem 6], and we assume a standard random access model of computation, with a uniform cost measure.
@@ 3.3 Algorithmic aspects
There exists an algorithm that, given a well-behaved SPJU plan $p$ of intermediate degree $\leq d$ and a structure $\mathbf{D}$, evaluates $p$ on $\mathbf{D}$ in time $O(|p|^{2}*|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d})$.
@@
To argue this proposition, we consider an algorithm that evaluates the plan $p$ inductively. It suffices to show that each subplan $q\in\mathsf{subplans}(p)$ can be computed in time $O(|p|*|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d})$. By the definition of intermediate degree, each subplan $q$ has at most $O(M_{\mathbf{D}}^{d})$ answers (tuples in $\mathsf{out}(p,\mathbf{D})$). When $q$ is a basic plan, the required bound holds immediately. When $q$ is a select or project subplan, i.e., of the form $\sigma_{\theta}(q^{\prime})$ or $\pi_{j_{1},\ldots,j_{n}}(q^{\prime})$, respectively, the answers of $q$ can be naturally derived from the answers of the immediate subplan $q^{\prime}$, within the given time bound. When $q$ is a union subplan, i.e., of the form $q_{1}\cup q_{2}$, it is immediate that the answers of $q$ can be derived by combining the answers of $q_{1}$ and $q_{2}$ and eliminating duplicates, which is feasible in the given time bound. The key point is that, when $q$ is a join subplan, by well-behavedness, it is of the form $\Join_{\theta}(q_{1},\ldots,q_{\ell})$ and $\mathsf{out}(q,\mathbf{D})$ can be computed by maintaining a relation that is a subset of $\mathsf{out}(q_{1},\mathbf{D})\times\mathsf{dom}(\mathbf{D})$. When evaluating the multiway join, we can thus evaluate the sequence of binary joins $\Join_{\theta_{\ell-1}}(\ldots(\Join_{\theta_{2}}(\Join_{\theta_{1}}(q_{1},q_{2}),q_{3}),\ldots),q_{\ell})$, where $\theta_{1},\ldots,\theta_{\ell-1}$ are the corresponding equalities from $\theta$ for the join partners in question. Each binary join can be evaluated in time $O(|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d})$ by [9], and in total, there are at most $|p|$ many joins, giving us the desired runtime of $O(|p|*|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d})$ to evaluate the subplan $q$.
4 SPJ Plans
@@
This section is dedicated to proving Theorem 3.4, which, apart from being of interest in its own right, serves as a building block towards proving our main result in the next section. Proving Theorem 3.4 essentially requires us to show two statements: (i) we can synthesize a plan, which is $\Sigma$-equivalent to the original one and has intermediate degree $\leq d$ and (ii) any plan which is $\Sigma$-equivalent to the original one has a subplan with output degree at least $d$. We will formalize and prove (i) and (ii) in Theorems 4.7 and 4.8, respectively, in Subsection 4.2. To show the theorem, we make two more basic observations on the introduced degree notions. We begin by presenting a definition that will aid us in arguing that a plan has a certain output degree. We say that $p$ has output degree $\geq d$ if there exists a function $g:\mathbb{N}\to\mathbb{N}$ with $g(M)\in\Omega(M^{d})$ and there exists an infinite sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures with $M_{\mathbf{D}_{n}}$ unbounded such that (for all $n\geq 1$) $|\mathsf{out}(p,\mathbf{D}_{n})|\geq g(M_{\mathbf{D}_{n}})$. We then have the following lemma.
@@
Theorem 3.4 can then be argued as follows: Given a plan $p$, we first compute the open structure $(\mathbf{A},\overline{a})$, which can be done in polynomial time by Theorem 4.6. Then, the algorithm of Theorem 3.4 computes $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, again in polynomial time. It then computes the core of this open structure; this can be performed in exponential time. Let $(\mathbf{C},\overline{c})$ be this core. Note that the size of $(\mathbf{C},\overline{c})$ is bounded above by the size of $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, since the core of a structure is a substructure of the structure. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. From $(\mathbf{C},\overline{c})$, we can then synthesize the desired plan $p^{\prime}$ of intermediate degree $\leq d$ by statement (i), formalized in Theorem 4.7, from above. Consider any plan $p^{\prime\prime}$ that is $\Sigma$-semantically equivalent to $p^{\prime}$. By statement (ii), formalized in Theorem 4.8, $p^{\prime\prime}$ has a subplan with output degree $\geq d$. For any $\epsilon>0$, by Lemma 4.1, this subplan does not have output degree $\leq(d-\epsilon)$, and so by Lemma 4.2, $p^{\prime\prime}$ does not have intermediate degree $\leq(d-\epsilon)$. Hence, we have computed a $\Sigma$-equivalent plan $p^{\prime}$ of polynomial size and of intermediate degree $d$, which is the best possible intermediate degree. In the following, we will formalize the connection from plans to open structures and auxiliary notions, which are needed to prove the theorems formalizing statements (i) and (ii) in Subsection 4.2.
4.1 Relating Queries to Structures
There is a known correspondence between Boolean SPJ plans and structures [5]; by a Boolean SPJ plan, we mean one that evaluates to true or false (or alternatively, a relation of arity 0) on a given structure $\mathbf{D}$. This correspondence yields that each such plan can be translated to a structure $\mathbf{A}$ such that, for an arbitrary structure $\mathbf{D}$, the plan is true on $\mathbf{D}$ if and only if there is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$. We prove a version of this correspondence in Theorem 4.6 for SPJ plans that can output relations of any finite arity, and not just relations of arity 0. In order to formulate our version, we utilize the notion of open structure.
@@ 4.1 Relating Queries to Structures
Apart from showing that any SPJ plan $p$ can be translated to a $p$-representation, we also want to provide a tree decomposition of this open structure with certain additional properties, which we call $p$-decomposition. This allows us to relate plans to structural, decomposition-based measures. While the translation that we present is, on a high level, along the lines of known translations from SPJ plans to conjunctive queries, we emphasize that our translation also provides a tree decomposition that will allow us to analyze the intermediate degree. In particular, for a given plan, each subplan will correspond to a vertex of the tree decomposition whose associated set of elements corresponds to the free variables of the subplan. We say a tree decomposition $(T,\chi)$ is rooted if there is a distinguished node $v_{0}\in V(T)$, called the root of $T$. The formal definition of $p$-decomposition follows.
@@
Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ a $p$-representation. A $p$-decomposition of $(\mathbf{A},\overline{a})$ is a rooted tree decomposition $E=(T,\chi)$ of $(\mathbf{A},\overline{a})$ such that there exists a bijection $\alpha:\mathsf{subplans}(p)\to V(T)$, where $\alpha(p)$ is the root of $T$, and a map $\beta:\mathsf{subplans}(p)\to\mathsf{dom}(\mathbf{A})^{*}$, where $\beta(p)$ is equal to $\overline{a}$, where for each $q\in\mathsf{subplans}(p)$, we have that $\{\beta(q)\}=\chi(\alpha(q))$.
In the definition of $p$-decomposition, each subplan $q$ of $p$ corresponds to a vertex of the tree $T$, via $\alpha$, and for each such subplan $q$, it holds that $\beta(q)$ is a tuple that lists, in some order (and possibly with repetition), the elements of the set to which $\chi$ maps q’s tree vertex. Let $p$ be an SPJ plan over signature $\sigma$, $(\mathbf{A},\overline{a})$ be a $p$-representation, and $E=(T,\chi)$ be a rooted $p$-decomposition $E=(T,\chi)$ of $(\mathbf{A},\overline{a})$ with associated mappings $\alpha$ and $\beta$. For a vertex $v\in T$, let $S_{v}$ be the union of $\chi(u)$ over all descendants $u$ of $v$ (including $v$ itself) and let $\mathbf{A}_{\leq v,E}$ denote the induced substructure $\mathbf{A}[S_{v}]$. We say a subplan $q\in\mathsf{subplans}(p)$ satisfies the containment property w.r.t. a $p$-decomposition if, for each structure $\mathbf{D}$ over $\sigma$, the relationship $\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}_{\leq\alpha(q),E},\beta(q),\mathbf{D})$ holds. Further, we say a $p$-decomposition satisfies the containment property, if every subplan $q\in\mathsf{subplans}(p)$ satisfies the containment property w.r.t. the $p$-decomposition. The containment property will allow us to lower bound the sizes of intermediate relations of $p$ using a $p$-representation along with a $p$-decomposition satisfying the containment property.
@@
Consider the signature $\sigma$ containing the binary relations $R/2$ and $S/2$ and the SPJ plan $p=\Join_{(1=3)}(R,\pi_{1}(S))$ over $\sigma$. Then it can be readily seen, with structure $\mathbf{A}=\{R(a,b),S(a,d)\}$ and tuple $\overline{a}=(a,b,a)$, that $(\mathbf{A},\overline{a})$ is a $p$-representation. Consider the following trees: [Uncaptioned image] On the left, we have the plan $p$ represented as a tree (each node represents the subplan of its subtree), while on the right we have some tree $T$. We provide mappings $\alpha$, $\beta$, and $\chi$ such that $(T,\chi)$ is a $p$-decomposition via mappings $\alpha$ and $\beta$. First, $\alpha$ is the obvious bijection from the subplans of $p$ to $V(T)$ with $\alpha(p)=v_{0}$, $\alpha(R)=v_{1}$, $\alpha(\pi_{1}(S))=v_{2}$, and $\alpha(S)=v_{3}$. Further, we define $\beta$ as $\beta(p)=\overline{a}$, $\beta(R)=(a,b)$, $\beta(\pi_{1}(S))=(a)$, and $\beta(S)=(a,d)$. Lastly, define $\chi(\alpha(q))$ as $\{\beta(q)\}$ for all subplans $q$ of $p$. It can be readily verified that $(T,\chi)$ is indeed a $p$-decomposition, and that it satisfies the containment property.
We show with the following theorem, that a $p$-representation and a $p$-decomposition satisfying the containment property always exist and are efficiently computable, via an inductive argument over the structure of $p$ given in the appendix.